# Stability of quasi-particle creation and multiband geometry in fractional Chern insulators under magnetic fields

### v1

@@ VI Summary
In this work, we have discussed the stability of the quasi-particle creation in fractional Chern insulators under magnetic fields $\phi$. We considered two models with distinct quantum geometric properties, namely the Kapit -Mueller model and the checkerboard model. We demonstrated by the exact diagonalization that the Kapit-Mueller model with $\phi\neq 0$ exhibits the stable quasi-particle creation. On the other hand, the FCI state in the checkerboard model is unstable under $\phi<0$ and the non-FCI state is realized even for moderate interactions, although the quasi-electron creation is stable. The model dependence in the stability of the quasi-particles can be attributed to the difference in the multiband quantum geometry under the magnetic field. Therefore, multiband geometric structures together with the multiband bandwidth are essential ingredients to host stable quasi-particle creation.
@@ Acknowledgements.
We are grateful to Nobuyuki Okuma and Manato Fujimoto for fruitful discussions. This work is supported by JSPS KAKENHI Grant No. 22K03513.
Appendix A Quasi-particle creation by increasing unit cells
@@
Here, we quickly discuss the quasi-particle creation in the checkerboard model at $\phi=0$ caused by a change of the system size, where the band properties are kept intact. We consider the system with $N=5,U=1.0$ and increase one unit cell as $N_{s}=30\to 32$. The spectral flow is shown in Fig.13. The counting rule at $V_{p}=0$ is #$\#_{\text{qh}}=16$ and they lead to the three fold quasi-degenerate ground state at $V_{p}=1$. The one-plaquette Chern number in the parameter space is smooth and flat as shown in Fig.13 (c), along with the many-body Chern number $C=-1$. These results suggest that finite size effects for the quasi-particle creation are negligible even for the small system size $N=5$.
In addition, we have also confirmed creation of the quasi-particles in the Kapit-Muller model with $N=4$ by increasing and decreasing one unit cell. It is found that the creation of the quasi-particles against magnetic fields is stable and the finite size effects are safely negligible.
@@ Appendix B Spectral flow in the checkerboard model with $N=4,5$
Figure 14: Spectral flow in the checkerboard model for the system under $N_{\phi}=-1$. The system sizes are (a)$\sim$(d)($N_{s}=30$,$N=5$,#$\#_{\text{qh}}=16$) and (e)$\sim$(h)($N_{s}=24$,$N=4$,#$\#{\text{qh}}=13$). The blue curves for $V_{p}=0$ are the states expected from the counting rule, and those for $V_{p}=10$ are the three lowest energy levels as in Fig.5.
We show the spectral flows for $N=4,5$ in $N_{\phi}=-1$ in the checkerboard model in Fig. 14. For $N=5$, the counting rule #$\#_{\text{qh}}=16$ does not hold at $U=0\sim 2.0$ because there is no clear gap between the low energy 16 states and other excited states (Fig. 14(a)). The three fold quasi-degeneracy also does not hold when the pinning potential $V_{p}$ is introduced in Fig. 14(b), which indicates that the quasi-hole cannot be created and a non-FCI state is formed. The phase transition to the FCI with the pinned quasi-hole occurs around $U=2.0\sim 3.0$. The counting-rule and the three fold quasi-degeneracy hold under $U\geq 3.0$ with the clear energy gap as seen in Fig 14(c),(d), which is consistent with the result of the variance of the one-plaquette Chern number. Similarly for $N=4$, there is an energy gap for $U=3.0\sim 8.0$ where there are #$\#_{\text{qh}}=11$ low energy states (Fig. 14(e)$\sim$(h)) and the variance $\mathcal{V}[\mathcal{C}]$ is small.
For $N_{\phi}=1$, energy gaps above low energy states open even at $U=1.0$ for both $N=4,5$ as shown in Fig. 15. The energy gap develops and the FCI with the pinned quasi-electron becomes more stable as the interaction increases. The stabilization of the FCI is seen also in the variance $\mathcal{V}[\mathcal{C}]$. We concluded that the quasi-electron creation is stable in a wide range of the interaction for $N=4,5$.

### v2

@@ VI Summary
In this work, we have discussed the stability of the quasiparticle creation in fractional Chern insulators under magnetic fields $\phi$. We considered two models with distinct quantum geometric properties, namely the Kapit -Mueller model and the checkerboard model. We demonstrated by the exact diagonalization that the Kapit-Mueller model with $\phi\neq 0$ exhibits the stable quasiparticle creation. On the other hand, the FCI state in the checkerboard model is unstable under $\phi<0$ and the non-FCI state is realized even for moderate interactions, although the quasielectron creation is stable. The model dependence in the stability of the quasiparticles can be attributed to the difference in the multiband quantum geometry under the magnetic field. Therefore, multiband geometric structures together with the multiband bandwidth are essential ingredients to host stable quasiparticle creation.
@@ Acknowledgements.
We are grateful to Emil J. Bergholtz, Nobuyuki Okuma, and Manato Fujimoto for fruitful discussions. This work is supported by JSPS KAKENHI Grant No. 22K03513.
Appendix A quasiparticle creation by increasing unit cells
@@
Here, we quickly discuss the quasiparticle creation in the checkerboard model at $\phi=0$ caused by a change of the system size, where the band properties are kept intact. We consider the system with $N=5,U=1.0$ and increase one unit cell as $N_{s}=30\to 32$. The spectral flow is shown in Fig.13. The counting rule at $V_{p}=0$ is #$\#_{\text{qh}}=16$ and they lead to the three fold quasi-degenerate ground state at $V_{p}=1$. The one-plaquette Chern number in the parameter space is smooth and flat as shown in Fig.13 (c), along with the many-body Chern number $C=-1$. These results suggest that finite size effects for the quasiparticle creation are negligible even for the small system size $N=5$.
In addition, we have also confirmed creation of the quasiparticles in the Kapit-Muller model with $N=4$ by increasing and decreasing one unit cell. It is found that the creation of the quasiparticles against magnetic fields is stable and the finite size effects are safely negligible.
@@ Appendix B Spectral flow in the checkerboard model with $N=4,5$
Figure 14: Spectral flow in the checkerboard model for the system under $N_{\phi}=-1$. The system sizes are (a)$\sim$(d)($N_{s}=30$,$N=5$,#$\#_{\text{qh}}=16$) and (e)$\sim$(h)($N_{s}=24$,$N=4$,#$\#_{\text{qh}}=13$). The blue curves for $V_{p}=0$ are the states expected from the counting rule, and those for $V_{p}=10$ are the three lowest energy levels as in Fig.5.
We show the spectral flows for $N=4,5$ in $N_{\phi}=-1$ in the checkerboard model in Fig. 14. For $N=5$, the counting rule #$\#_{\text{qh}}=16$ does not hold at $U=0\sim 2.0$ because there is no clear gap between the low energy 16 states and other excited states (Fig. 14(a)). The three fold quasi-degeneracy also does not hold when the pinning potential $V_{p}$ is introduced in Fig. 14(b), which indicates that the quasihole cannot be created and a non-FCI state is formed. The phase transition to the FCI with the pinned quasihole occurs around $U=2.0\sim 3.0$. The counting-rule and the three fold quasi-degeneracy hold under $U\geq 3.0$ with the clear energy gap as seen in Fig 14(c),(d), which is consistent with the result of the variance of the one-plaquette Chern number. Similarly for $N=4$, energy gaps above low energy states develop between $U=3.0\sim 8.0$ (Fig. 14(e)$\sim$(h)) and the variance $\mathcal{V}[\mathcal{C}]$ is small for $U\geq 5.0$.
For $N_{\phi}=1$, energy gaps above low energy states open even at $U=1.0$ for both $N=4,5$ as shown in Fig. 15. The energy gap develops and the FCI with the pinned quasielectron becomes more stable as the interaction increases. The stabilization of the FCI is seen also in the variance $\mathcal{V}[\mathcal{C}]$. We concluded that the quasielectron creation is stable in a wide range of the interaction for $N=4,5$. Note that the stability of the quasielectron creation only weakly depends on the strength of the pinning potential, and we have confirmed that the results are basically unchanged for $V_{p}=-0.1,-1,-10$. Essentially same results can also be obtained without the projection.